# Data Quality Matters: Suicide Intention Detection on Social Media Posts Using a RoBERTa-CNN Model

### v1

@@ V Discussion
This paper uses the RoBERTa-CNN model to do SID because of its effectiveness in language tasks. The results in Section IV-D show its good performance on text data. To be specific, the attention mechanism helps RoBERTa be more constructive in capturing contextual information. CNN reduces the limitation of the RoBERTa model by capturing the local features with its Conv layer and identifies insight linguistic patterns in a sentence [33]. Thus, the CNN head enhances the RoBERTa-CNN’s robustness and accuracy.
Our integration structure leverages the strength of both RoBERTa and CNN for SID, which enables RoBERTa-CNN to surpass the results of CNN-BiLSTM and XGBoost [32] on SDD dataset.
Simultaneously, the RoBERTa-CNN also has limitations. Due to its complexity, the RoBERTa-CNN requires huge computational costs as the Max_Len increases. To address this issue, we will research dimension deduction [34]. The simple structure of the CNN head drives us to find more power classifiers such as XnODR/XnIDR [35] and Multi-channel classifier [36] too. Finally, there is very few GPT-related research on SID, which motivates us to use the GPT model to do contrastive experiments on the SID task.
@@ VI Conclusions
This paper presented a study on SID leveraging RoBERTa-CNN. We tested the model on the SDD dataset. The promising results state the robustness of the model in identifying the textual signs of suicide ideation. Moreover, RoBERTa-CNN outperforms several existing models in the same dataset. This convinces us that RoBERTa-CNN does SID well on text data. Future work also includes developing a multi-modal model to do SID by analyzing visual, text, and speech data.
@@ Acknowledgment
We would like to thank Miss. Zhuo Chen who helped us clean the data and Mr. Jiahan Li who joined our weekly meetings and contributed to our discussions.

### v2

@@ V Discussion
This paper uses RoBERTa-CNN to conduct SID because of its effectiveness in language tasks. The results in Section IV-D show its power on text data. The RoBERTa model with its attention mechanism is powerful in capturing contextual information. CNN reduces the limitation of the RoBERTa model by capturing the local features with its Conv layer and identifies insight linguistic patterns in a sentence [27]. Thus, the CNN head enhances the robustness and accuracy of RoBERTa-CNN.
Moreover, the data cleaning removes the noise from the original SDD dataset and improves the data quality, substantially contributing to RoBERTa-CNN’s solid performance.
Nevertheless, RoBERTa-CNN has some limitations. Due to its complexity, RoBERTa-CNN requires huge computational costs as the Max_Len increases. Light-weight RoBERTa may solve the problem. The simple structure of the CNN head drives us to find more power classifiers such as XnODR/XnIDR [28] and Multi-channel classifier [29]. Finally, there is very little ChatGPT-related research on SID now, which motivates us to use the ChatGPT model to do contrastive experiments on the SID task.
@@ VI Conclusions
This paper presented a study on SID leveraging RoBERTa-CNN. We tested the model on the SDD dataset. The promising results imply the robustness of the model in identifying the textual signs of suicide ideation, convincing us that RoBERTa-CNN does SID well on text data. Future work includes developing a multi-modal model to do SID by analyzing visual, text, and speech data.
@@ Acknowledgment
The author would like to thank Miss. Zhuo Chen for helping clean the data, and Mr. Jiahan Li for joining weekly meetings and contributing to the discussions. The research was partially supported by a Partners in Scholarship (PINs) summer research support to Emily Lin.